OpenStudy (anonymous):

how do you do this ? Write the slope intercept for the line that passes through (2,-2) and (-1,10)

OpenStudy (anonymous):

Slope intercept form is given by: \[\large y = mx + c\] Where m is the slope and c is y intercept.. Now use: \[\large \frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1} \] Are you able to do this??

OpenStudy (anonymous):

y2-y1 /x2-x1 y=-4x+6 -2 - 10 / 2 +1 y=-4x+14 -12 / 3 -4=slope y=-4x + b -2 = -4(2)+b -2 = -8 +b +8 +8 6 = b y=-4x+b 10=4(-1)+b 10=-4+b +4 +4 14 = b this is what i got did i do it right

OpenStudy (anonymous):

Let me check..

OpenStudy (anonymous):

Did you find b for both???

OpenStudy (anonymous):

I got 6 and 14 for b

OpenStudy (anonymous):

I don't know what we do when b has two values.. Let me solve this as I have learned: \[\large \frac{y +2}{1 + 2} = \frac{x - 2}{-1 -2}\] \[\large \frac{y+2}{12} =\frac{x - 2}{-3} \implies -3y - 6 = 12x - 24\] \[\large -3y = 12x - 18 \implies 3y = -12x + 18 \implies y = -4x + 6\]

OpenStudy (anonymous):

it is not 1 + 2 it is 10 + 2 in the denominator in LHS.. Sorry there is my mistake..

OpenStudy (anonymous):

I found your mistake @pige77 y=-4x+b 10=-4(-1)+b 10= 4+b -4 -4 6 = b You will get 6 both this also..

OpenStudy (anonymous):

Are you getting it ??? @pige77

OpenStudy (anonymous):

How you made -4 4 there?? y = -4x + b y is 10 and x is -1 So: 10 = -4(-1) + b Right ??

OpenStudy (anonymous):

yes thank you it is those little mistakes thae make a huge error

OpenStudy (anonymous):

My English Oh My God..

OpenStudy (anonymous):

will you help with another slope intercept equation

OpenStudy (anonymous):

Yeah sure...

OpenStudy (anonymous):

ok give me a sec not good typing fast

OpenStudy (anonymous):

Write the slope intercept equation for the line with slope -2/3 and y-intercept (0,3)

OpenStudy (anonymous):

What do you think here b is ??

OpenStudy (anonymous):

i got y=-2/3x+3

OpenStudy (anonymous):

Well Done...

OpenStudy (anonymous):

ty one more

OpenStudy (anonymous):

Go on..

OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (-2,3) and is perpendicular to 2x-9y=-12.

OpenStudy (anonymous):

i got y=9/2+12 i get my signs mixed up

OpenStudy (anonymous):

Slope will be -9/2 I guess..

OpenStudy (anonymous):

2x-9y=-12 -9y = -12 - 2x 9y = 2x + 12 y = (2/9)x + 12/9 Now what is negative reciprocal of 2/9 ??

OpenStudy (anonymous):

-9/2

OpenStudy (anonymous):

so where did i go wrong?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Let me check whole..

OpenStudy (anonymous):

y-y1=m(x-x1) y-3=9/2(x-(-2)) y-3=9/2x+9 +3 +3 y=9/2+12

OpenStudy (anonymous):

Use equation: \[\large y - y_1 = m(x - x_1)\] y - 3 = -9/2 (x + 2) Now: y - 3 = -9/2x - 9 y = -9/2x - 6

OpenStudy (anonymous):

Why are you using 9/2 ??? Slope is -9/2 no??

OpenStudy (anonymous):

did the signs mess me up then

OpenStudy (anonymous):

so the answer is y=-9/2x-6 right

OpenStudy (anonymous):

Yes it is right//

OpenStudy (anonymous):

ty sooooooo much

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

you have been a huge help now i have to figure out how to graph some of these

OpenStudy (anonymous):

I am not good at graphing.. I will see it what is the question..

OpenStudy (anonymous):

the one that we did that the answer is y=-2/3+3 is one of them

OpenStudy (anonymous):

i dont know how to try them on here

OpenStudy (anonymous):

to show you

OpenStudy (anonymous):

See you have y intercept here 3.. and slope is negative 2/3 so the line will be from upper left to lower right on the graph..  Somewhat like this...